# Contextualized Messages Boost Graph Representations

### v2

@@ Appendix A Proofs
Definition 2 (Pseudometric).
Let $\mathcal{H}$ be a non-empty set. A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$.
  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
Definition 3 (Conditionally positive definite kernel [36]).
@@
with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and
$\sum_{i=1}^{N}c_{i}=0.$ (22)
Lemma 1.
The negative of a pseudometric on $\mathcal{H}$ is a conditionally positive definite kernel on $\mathcal{H}$.
Proof.
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric. For all $N\in\mathbb{N}$ and $\bm{h}^{(0)},\bm{h}^{(1)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$, by the triangle inequality of $d$,
$-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)$ $\geq-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)$ (23)
$=-\sum_{j=1}^{N}c_{j}\sum_{i=1}^{N}c_{i}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}c_{i}\sum_{j=1}^{N}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)=0.$ (24)
Theorem 2 (Hilbert space representation of conditionally positive definite kernels [3, 35, 36]).
@@
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric. From Lemma 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
$d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}.$ (26)
For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}$ (27)
$\delta_{2}^{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}<\delta_{1}^{2}$ (28)
$\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (29)
@@
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (31)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$. From Lemma 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
@@
Note that $D$ is indeed a pseudometric since $k$ is positive definite as noted by Joshi et al. [23]. [^2] If $k$ is also integrally strictly positive definite [37], then the hash function $G$ becomes injective. By the linearity of the inner product, it may be shown that
$D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert,$ (33)
@@
and $g$ is the corresponding feature map of the kernel $k$. For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
$\delta_{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert<\delta_{1}$ (35)
$\delta_{2}^{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert^{2}<\delta_{1}^{2}$ (36)
$\varepsilon_{2}<D\left(\bm{H_{1}},\bm{H_{2}}\right)<\varepsilon_{1}.$ (37)
@@ Appendix B Experimental set-up
All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (dgl, version 2.1.0+cu121, Apache License 2.0) [43] with PyTorch (torch, version 2.2.1, BSD-3) [33] backend. For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values. For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials. The hyperparameters are chosen based on previous results and heuristics without extensive tuning. The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
@@ B.1 Synthetic datasets
Adopting Brody et al. [7], the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration. All models utilize a single GNN layer with $4n$ hidden units. A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN. Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
@@
The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with randomly assigned labels from one of $c$ classes, while the test dataset comprises 1,000 directed graphs with the same configuration. All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function. A feed-forward neural network is also used for GIN. Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
@@ B.2 Benchmark datasets
The dataset is obtained from dgl (version 2.1.0+cu121, Apache License 2.0) with data splits (training, validation, test) following Dwivedi et al. [12]. In line with Dwivedi et al. [12], all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000. Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting. The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al. [41]. Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10. Model training is performed with the AdamW [29] optimizer for over a maximum of 500 epochs with a batch size of 128 (whenever applicable) and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
@@
The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al. [19]. The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections. Regularization with weight $1\times 10^{-6}$ and dropouts with rates in increments of 0.1 are also used to prevent overfitting. The symmetric mean aggregator is used along with existing tricks in literature. Model training is performed with the AdamW [29] optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss.
@@
The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al. [19]. The models utilize a single GNN layer, modified to leverage edge features as described in Appendix C, with 300 hidden units, batch/graph normalization, and residual connections. Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting. The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function. Model training is performed with the AdamW [29] optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss.
Appendix C SIR-GCN extensions
@@
$\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (39)
with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA. Similarly, this extension may be viewed as a generalization of GIN with edge features [18].
Furthermore, one may inject inductive bias into the pseudometric $D$ which may correspond to specifying the architecture type for the corresponding message function $g$. For instance, if node features are known to have a sequential relationship (e.g. stock [17] and fMRI [24] data), $g$ may then be aptly modeled using recurrent-type networks.

### v3

@@ Appendix A Proofs
Definition 2 (Conditionally positive definite kernel (Schölkopf, 2000)).
@@
with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and $\sum_{i=1}^{N}c_{i}=0$.
Theorem 2 (Hilbert space representation of conditionally positive definite kernels (Berg et al., 1984; Schoenberg, 1938; Schölkopf, 2000)).
@@
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric. From Assumption 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert=d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (23)
Hence, for every $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (24)
@@
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (26)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$. From Assumption 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
@@
Note that $D$ is indeed a pseudometric since $k$ is positive definite as noted by Joshi et al. (2011). [^3] If $k$ is also integrally strictly positive definite (Sriperumbudur et al., 2010), then the hash function $G$ becomes injective and $D$ becomes a metric. By the reproducing property of $k$ and the linearity of the inner product, it may be shown that
$\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert=D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right),$ (28)
@@
and $g$ is the corresponding feature map of the kernel $k$. Hence, for every $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1}.$ (30)
@@ Appendix B Experimental set-up
All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (Wang et al., 2019a) with PyTorch (Paszke et al., 2019) backend. For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values. For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials. The hyperparameters are chosen based on previous results and heuristics without extensive tuning. The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
@@ B.1 Synthetic datasets
Adopting Brody et al. (2021), the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration. All models utilize a single GNN layer with $4n$ hidden units. A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN while PNA uses the sum, max, and std aggregators. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
@@
The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with randomly assigned labels from one of $c$ classes, while the test dataset comprises 1,000 directed graphs with the same configuration. All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function. A feed-forward neural network is also used for GIN while PNA uses the sum, max, and std aggregators. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
@@ B.2 Benchmark datasets
The datasets are obtained from dgl with data splits (training, validation, test) following Dwivedi et al. (2023). In line with Dwivedi et al. (2023), all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000. Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting. The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al. (2019). Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over a maximum of 500 epochs with a batch size of 128, whenever applicable, and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss. The reported results for other models in Table 3 are obtained from Dwivedi et al. (2023), Corso et al. (2020), and Tailor et al. (2021).
@@
The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al. (2020). The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections. Regularization with weight $1\times 10^{-6}$ and dropouts with rates in increments of 0.1 are also used to prevent overfitting. The symmetric mean aggregator is used along with existing tricks in literature. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss. The reported results for other models in Table 4 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
@@
The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al. (2020). The models utilize a single GNN layer, modified to leverage edge features as described in Appendix D, with 300 hidden units, batch/graph normalization, and residual connections. Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting. The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss. The reported results for other models in Table 5 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
Appendix C Model run time
As an additional evaluation, the run time per epoch for each model in the synthetic datasets is presented in Tables 6 and 7. The results, when considered alongside Tables 1 and 2, illustrate that SIR-GCN achieves a balance between computational complexity and model expressivity. Notably, SIR-GCN is computationally less expensive than PNA while delivering superior performance. These results further underscore the practicality of the proposed model.
Table 6: Run time per epoch on DictionaryLookup.
Model $n=10$ $n=20$ $n=30$ $n=40$ $n=50$
GCN 0.3526s $\pm$ 0.0778s 0.4734s $\pm$ 0.0468s 0.4777s $\pm$ 0.0854s 0.5619s $\pm$ 0.0518s 0.5520s $\pm$ 0.0679s
GraphSAGE 0.4565s $\pm$ 0.0873s 0.5264s $\pm$ 0.0317s 0.5716s $\pm$ 0.1132s 0.7742s $\pm$ 0.0597s 0.9193s $\pm$ 0.0473s
GATv2 0.3950s $\pm$ 0.1017s 0.5276s $\pm$ 0.0556s 0.6191s $\pm$ 0.0879s 0.7472s $\pm$ 0.0346s 1.0065s $\pm$ 0.0280s
GIN 0.3696s $\pm$ 0.0899s 0.4610s $\pm$ 0.0459s 0.4670s $\pm$ 0.0781s 0.5947s $\pm$ 0.0548s 0.5194s $\pm$ 0.0993s
PNA 0.8854s $\pm$ 0.0412s 1.1913s $\pm$ 0.1024s 1.4526s $\pm$ 0.0684s 1.8793s $\pm$ 0.0528s 2.8387s $\pm$ 0.0603s
SIR-GCN 0.4687s $\pm$ 0.0777s 0.6066s $\pm$ 0.0398s 0.8053s $\pm$ 0.0485s 1.1496s $\pm$ 0.0427s 1.7031s $\pm$ 0.0458s
Table 7: Run time per epoch on GraphHeterophily.
Model $c=2$ $c=4$ $c=6$ $c=8$ $c=10$
GCN 0.4243s $\pm$ 0.0520s 0.3852s $\pm$ 0.0517s 0.3868s $\pm$ 0.0743s 0.4166s $\pm$ 0.0551s 0.4177s $\pm$ 0.0494s
GraphSAGE 0.4691s $\pm$ 0.0400s 0.4790s $\pm$ 0.0440s 0.4399s $\pm$ 0.0629s 0.4501s $\pm$ 0.0603s 0.4964s $\pm$ 0.0601s
GATv2 0.4710s $\pm$ 0.0978s 0.4941s $\pm$ 0.0567s 0.4718s $\pm$ 0.0361s 0.5514s $\pm$ 0.0608s 0.5437s $\pm$ 0.0724s
GIN 0.4085s $\pm$ 0.0741s 0.3875s $\pm$ 0.0627s 0.3855s $\pm$ 0.0645s 0.4298s $\pm$ 0.0566s 0.4329s $\pm$ 0.0534s
PNA 2.2963s $\pm$ 0.0413s 2.4238s $\pm$ 0.0611s 2.4577s $\pm$ 0.0533s 2.4741s $\pm$ 0.0665s 2.5623s $\pm$ 0.0425s
SIR-GCN 0.5338s $\pm$ 0.0353s 0.5264s $\pm$ 0.0737s 0.5635s $\pm$ 0.0695s 0.5764s $\pm$ 0.0401s 0.6230s $\pm$ 0.0388s
Appendix D SIR-GCN extensions
@@
$\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (32)
with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA. Similarly, this extension may be viewed as a generalization of GIN with edge features (Hu et al., 2019).
Furthermore, one may inject inductive bias into the pseudometrics which may correspond to specifying the architecture type for the corresponding message function $g$. For instance, if node features are known to have a sequential relationship (e.g., stock (Hsu et al., 2021) and fMRI (Kim & Ye, 2020) data), $g$ may then be aptly modeled using recurrent or convolutional networks.